

********************************************************************
UARS WEEKLY STATUS REPORT - 01 JAN 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
Happy New Year!
***********
SUSIM
***********
SUSIM was off between 12/23/98 and 12/30/98. On all other days SUSIM
successfully gathered its normal solar, offset, and occulation
data without apparent problems.
Reprocessing of all SUSIM experiment days to version 20 continues.
***********
PEM
***********
A Happy New Year to all from the PEM team!!
PEM conducting tests of the OBC activation. PEM activation occurs during
daylight only, when there is adequate power generated from the UARS solar
array. Currently, the PEM magnetometer and medium energy spectrometers
(particles below 30 keV) activate correctly under OBC control. Tests conducted
on the high energy particle spectrometers (HEPS -- particles greater than
30 keV) have had limited success. New OBC activation procedure for HEPS based
on previous tests has been created. Tests of the new activation sequence has
begun. The first of nine tests was successful. Due to testing, HEPS science
data from December 1 to December 15 and after December 17 should be considered
invalid and not used.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF has received a total of 452 CD-ROM requests as of 1-January-1999.
There were three orders received and shipped.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
70 days HAL (4.80) PROD (L3) (1874R-1721C-1F=152L)
31 days SOL (4.62) PROD (L2-L3) (2313R-2307C=6L)
132 days SUS (7.00) PROD (L3) (2540R-1100C-6F=1434L)
Plans continue for SOLSTICE to be run with debug in UCSS production. This
should aid in defining the cause of SOLSTICE wall clock time increases.
DAAC data created in the month of October 1998 for HALOE, HRDI, PEM, NMC,
and UKMO has been loaded into the DAAC Transfer Tables.
CDHF pending software upgrades: IDL 5.03, OpenIngres, UCSS 10.1, MSS 4.1,
NRT 2.5, UOAS 5.4 .
***********
FOT
***********
REPORTING PERIOD
This report covers 19 December 1998 (Orbit 39732, GMT Day 353) through 25
December 1998 (Orbit 39836, GMT Day 359).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) with
a +0.5 degree Roll (cold side down) and is performing nominally under two-
battery operations. The instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
WINDII 353/0005 39732 On
WINDII 353/0144 39733 Off
WINDII 354/0302 39749 On
ACRIM 357/2209 39806 Off
SUSIM 357/2209 39806 Off
PEM 357/2346 39807 Axis Off
WINDII 357/2346 39807 Off
MLS 358/0134 39808 On
The beta angle ranged between 16.7 degrees to 32.2 degrees for this report
period. The beta angle is now increasing toward a maximum angle of 34.0
degrees on 27 December 1998 (DOY 361).
As beta angle (and average power margin) increased this week, Battery 3
differential voltage decreased as expected. As a precaution, online and
offline personnel are monitoring the power subsystem with increased
vigilance for any unexpected behavior.
A +0.5 Degree Roll Offset Maneuver (cold side down) was performed on 17
December (DOY 351) in Orbit 39713 at 351/14:54:57. The maneuver was
requested by the WINDII instrument to facilitate the observation of Polar
Mesospheric Clouds. UARS will return to level flight on 11 January 1999
(DOY 011). The exact time for the return maneuver has not been finalized.
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.9 to 1.3 degrees C. Additional battery performance data
for each battery on 18 December 1998 and 25 December 1998 are provided in
Appendix D.
The clock error ranged from -11.3 to 11.4 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 01
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 03
% ACRIM 01
% MLS 10
% PEM 01
% SUSIM 01
% WINDII 04
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 03
Two (2) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/286 - 2287.496936
Transponder B: 98/300 - 2287.498964
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly
Reports are included in Appendix E. There are two open Anomaly Report.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/358 show the observatory data loss to be
21 hours, 40 minutes, 28.632 seconds (an increase of 01.216 seconds since
last report period). This is a 0.0341 percent data loss which equals a
99.9659 percent data capture for the mission.
The increase of 01.216 seconds occurred over two (2) days:
7 0.224 seconds on Day 356 due dat hit.
7 0.992 seconds on Day 357 due to hit in data (IP).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
Development of the procedure for the clock change to include the 01 January
1999 leap second adjustment is complete. A new ephemeris will be started
immediately after the OBC clock is adjusted. No impact on spacecraft
pointing is expected.
PEM testing, contingency operation simulation, and orbit adjust simulations
were suspended this week in order to support current operations. PEM
testing will resume on the 28th.
The FOT is investigating the feasibility of a new PMON action to change
battery V/T curves based on loading to mitigate battery risks in sun
pointing safehold.
The FOT discussed planning and other procedures for day-only operations and
mode transitions with instrument teams. SOLSTICE ON to DAY-ONLY transition
was successfully executed by ATC commands built from an instrument DAP on
day 358. It appears that since SOLSTICE is cycling they are not generating
their normal amount of heat and therefore they, along with the other SSPP
instruments, are getting a little cold. The FOT will analize the data for
this weekend and come up with some possible solutions for this problem.
YEAR 2000 OPERABILITY SUPPORT
UARS Test and Training Simulator (UTTS) - No change.
User Planning System (UPS) - No Change
Multi Satellite Operations Control Center (MSOCC) - No change.
Command Management System (CMS) - No Change
Flight Dynamics Facility (FDF) - No change.
UARS PROPULSION MODULE B/D THRUSTER LINE FILL
Planning continues for the simulator testing.
UARS RE-ENGINEERING - TELEMETRY PROCESSING
No change. A new release is expected in the near future.
MEETINGS
None
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is scheduled for scheduled for 11 January
1999.
The United States Naval Observatory (USNO) issued a Time Service
announcement advising that the International Earth Rotation Service (IERS)
will retard UTC time by 1.0 second at the end of December 1998. This means
the addition of one leap second as the last second of 1998.
The next annular solar eclipse will occur on 16 February 1999 over the
Indian Ocean, Australia and the South Pacific Ocean.
PERSONNEL CHANGES
none
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 18 December 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
39732 353/0005 98-503 WINDII Power ON
39733 353/0144 98-504 WINDII Pwr OFF
39749 354/0302 98-503 WINDII Power ON
39750 354/0446 98-505 OBC Clk Adj fr 20.6 to 0.0 ms/day
39806 357/2209 98-506 ACRIM Power OFF
39806 357/2209 98-507 SUSIM Power OFF
39807 357/2346 98-508 PEM AXIS Power Off
39807 357/2346 98-509 WINDII Power OFF
39808 358/0134 98-510 MLS Power ON
39810 358/0500 98-511 OBC Clk Adj fr 0.0 to 20.6 ms/day
39823 359/0139 98-512 MLS Excep Clear
39824 359/0322 98-512 MLS Excep Clear
39825 359/0503 98-512 MLS Excep Clear
39826 359/0643 98-512 MLS Excep Clear
39827 359/0823 98-513 MLS Scan Actuator OFF
39834 359/1853 98-514 MLS Scan Actuator ON
39835 359/2025 98-512 MLS Excep Clear
39835 359/2107 98-512 MLS Excep Clear
39836 359/2247 98-512 MLS Excep Clear
39836 359/2247 98-515 OBC Clk Adj fr 20.6 to 0.0 ms/day
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 18 December 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPSEQBIMON Y-HI PSE Quiet Bus Current Monitor 39769
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 25 Dec 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
39750 354/04:46 (20 Dec 98) from 20.6 to 00.0 msec/day
39810 358/05:04 (24 Dec 98) from 00.0 to 20.6 msec/day
39836 359/22:51 (25 Dec 98) from 20.6 to 00.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 25 Dec 1998
18 Dec 1998 - GMT Day 352 Beta = 16.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +67.2/+44.8 -3.05/-4.31 +0.0/ +0.0 16.8 0.000 0.0
2 +22.4/-44.8 +6.11/+4.83 +32.0/-14.4 25.3 1.024 21.9
3 +190.4/+22.4 +5.15/+3.87 +31.2/-15.2 25.1 1.022 21.7
25 Dec 1998 - GMT Day 3359 Beta = 32.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +61.6/+39.2 -2.73/-3.99 +0.0/ +0.0 16.8 0.000 0.0
2 +11.2/-67.2 +6.11/+5.15 +30.4/-15.2 25.3 1.028 21.8
3 +134.4/11.2 +5.15/+3.87 +30.0/-16.4 25.1 1.024 21.7
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 18 December 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-117 FOT 39737 353/0702 14170 Generic Late Acq
98-123 FOT 39788 356/1745 N/A AP/ODN DOWN
98-117 FOT 39806 357/2209 14170 Generic Late Acq
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-122 FOT 39639 346/18:14 CMD BUFFER PROB
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-117 FOT 39737 353/0702 14170 Generic Late Acq
98-123 FOT 39788 356/1745 N/A AP/ODN DOWN
98-117 FOT 39806 357/2209 14170 Generic Late Acq
SEVNTFW52